Find three consecutive positive integers such that the product of the first two is 22 less than 11 times the third
Answer by stanbon(75887)

Find three consecutive positive integers such that the product of the first two is 22 less than 11 times the third
------------------
1st: x-1
2nd: x
3rd: x+1
-----
Equation:
(x-1)x = 11(x+1)-22
------
x^2-x = 11x-11
----
x^2 - 12x + 11 = 0
(x-11)(x-1) = 0
---
x = 11 or x = 1
Positive solution:
x-1 = 10
x = 11
x+1 = 12
--------------------
Cheers,
Stan H.
---------------